OpenStudy (anonymous):

Help!! I don't know how to do this The endpoints of AB are A(2, 3) and B(8, 1). The perpendicular bisector of AB is CD, and point C lies on AB. The length of CD is (root of 10) units. The coordinates of point C are (-6, 2) (5, 2) (6, -2) (10, 4) . The slope of is -3 -1/3 1/3 3 . The possible coordinates of point D are (4, 5) (5, 5) (6, 5) (8, 3) and (2, 1) (4, -1) (5, -1) (6, -1) . Please pick 1 answer for all the questions

OpenStudy (anonymous):

Could you tell me what you're exactly looking for? The coordinate of C and the slope of the perpendicular bisector?

OpenStudy (anonymous):

@pooohbear99

OpenStudy (anonymous):

@yavar The coordinates of point C,The slope of CD,& The possible coordinates of point D(need 2)

OpenStudy (er.mohd.amir):

since CD is perpendicular bisector one point is mid point of A and B

OpenStudy (er.mohd.amir):

mid point is ((x1+x2)/2,(y1+y2)/2)

OpenStudy (dayakar):



OpenStudy (dayakar):

!) A(2,3) ,B(8,3) COORDINATES OF POINT C \[C =(\frac{ x1+x2 }{ 2 },\frac{ y1+y2 }{ 2 })\]

OpenStudy (dayakar):

\[C =(\frac{ 2+8 }{ 2 },\frac{ 3+3 }{ 2 })\] find C=

OpenStudy (dayakar):

C =(5.3)

OpenStudy (dayakar):

i mistaken B(8,1) then C = (5,2)

OpenStudy (dayakar):

if u are following i can continue